# CardioAI: A Multimodal AI-based System to Support Symptom Monitoring and Risk Detection of Cancer Treatment-Induced Cardiotoxicity

### v2

@@ 3. Participatory Design
We conduct a participatory design (PD) study with clinical experts in cancer treatment and treatment-induced cardiotoxicity for two primary purposes: 1) to better understand clinicians’ needs and challenges in the decision-making process of cardiotoxicity for cancer patients, 2) to derive design considerations for our proposed system, which aims to facilitate remote monitoring of multimodal patient data and early detection of cardiotoxicity with AI-CDSS support. Our system design will be grounded in the practical needs of clinicians’ current practice and focus on seamless integration into clinicians’ workflow without adding additional burdens, following the HCAI design principles (Amershi et al., 2019), which have been advocated and applied in prior work (Plana et al., 2014; Zhang et al., 2024; Albahri et al., 2023).
@@
3.1. Study Participants and Procedure
We recruited eleven clinicians through convenience sampling (Sedgwick, 2013) by reaching out to colleagues and professional connections in related fields. The clinical expertise of the participants includes both oncology and cardiology, as these are the two main roles in decision-making about cancer treatment-induced cardiotoxicity. Table 1 presents the demographics of the participants with a randomly assigned number denoted as $P\#$.
Before PD sessions, we derive an initial user interface (UI) design draft for a clinician-oriented information dashboard based on design considerations from the prior needs finding study (Wu et al., 2024), as shown in Figure 2. The initial UI design comprises conceptual drafts of multimodal data collection visualizations, such as vitals collected through wearables, symptoms collected via LLM-based voice assistants, AI-generated cardiotoxicity risk scores, and free-form text summaries of patient current health status. We use this design draft as a probe to understand clinicians’ feedback with respect to how AI technologies could help them in their clinical practices.
Our PD sessions consist of the following parts: (i) we first ask participants to describe their challenges in monitoring and detecting treatment-induced cardiotoxicity during their current clinical practice, and their perspectives on the potential of AI technologies to address these challenges; (ii) we then introduced, if needed, a brief overview of current technologies that have been used in RPM and AI-CDSS, such as wearables, LLM-based CA, and AI-driven prediction models; (iii) we present our UI design draft and collect feedback on visual design, interactive functionalities, and system specifications.
Figure 2. The initial UI design draft for clinician-oriented information dashboard
@@ 3.1. Study Participants and Procedure
The first author conducted a semi-structured interview with each participant online via Zoom. Each interview, which lasted 35 to 50 minutes, was recorded and transcribed into text transcriptions after the session. Two members of the research team independently coded the transcripts using inductive coding and thematic analysis (Braun and Clarke, 2012, 2019). The codes and themes were reviewed and refined into a consensus codebook through iterative discussions to ensure an accurate representation of the data by eliminating redundancy and merging overlapping codes. Both team members coded the transcripts for a second round after agreeing on the codebook, and we present the findings below.
@@ 3.2.1. Challenges in Clinical Decision-making for Cancer Treatment-Induced Cardiotoxicity
Previous studies (Cardinale et al., 2010; Kalam and Marwick, 2013, 2013; Wittayanukorn et al., 2017; Mascolo et al., 2021; Kamphuis et al., 2019; Pai and Nahata, 2000; Sheppard et al., 2013; Wu et al., 2024) have highlighted the unpredictable nature of treatment-induced cardiotoxicity, as symptoms can emerge at any time during or after treatment. Clinicians often rely on patients to self-report symptoms outside of clinical settings, however , patients may underreport or overlook symptoms due to health literacy barriers or concerns about bothering their healthcare providers. Our participants reinforced these known challenges, while also providing more nuanced insights based on their lived experiences.
Our participants reiterated the challenge of  subtle symptoms in the early stages in detecting treatment-induced cardiotoxicity, where such symptoms are difficult for patients to detect during routine clinical assessments or when patients are outside of clinical settings. Both P6 and P4 further emphasized the complexity by pointing out that symptoms may not only be subtle, but, in some cases, entirely absent even when patients are experiencing severe conditions. P4 shared a case where a patient, completely asymptomatic at home, was found to have a severe severe condition only because he was wearing a heart monitor, which is not provided routinely to all patients:
@@
Participants highlight several reasons beyond the lack of health literacy that may prevent patients from accurately reporting their symptoms. They noted logistical barriers that possibly hinder the effectiveness of self-reporting, even when patients are aware of their symptoms. These barriers, such as delayed communication between patients and healthcare providers, are particularly pronounced in resource-constrained settings, leading to discouragement and increased cognitive load on patients, as described by P1:
@@
Another limitation identified by the participants is the difficulty in recognizing cardiotoxicity-specific symptoms within current self-reporting systems. Many of the existing tools for symptom monitoring are designed for general purposes and do not focus on cardiotoxicity, which can lead to missed or underreported cardiac problems. For example, P5 emphasized that the surveys they use are “not specific to cardiotoxicity… unless there’s anything on the exam that suggests a cardiac problem.”
@@
Several participants (P1, P2, P7) expressed their frustration with the absence of reliable patterns or correlations between patient characteristics and cardiotoxicity risk, making it challenging to determine which patients require more intensive monitoring. As P1 noted, “We don’t have a good grasp on what features correlate with the likelihood of someone developing cardiotoxicity.” Such uncertainty often results in reactive decision-making, where clinicians have to wait for symptoms to manifest before intervening, rather than being able to anticipate and address risks sooner. Such delays, compounded by late symptom reporting and diagnostic processes, can lead to belated intervention. P2 highlighted this issue by sharing a case in which delays in reporting and testing prolonged treatment decisions: “The symptoms were reported late… everything was delayed, from the echo to the stress test to the EKG.
3.2.2. Opportunities for a Multimodal AI-based System for Monitoring and Early Detection of Treatment-Induced Cardiotoxicity
After we introduced the initial UI design draft of the information dashboard, participants expressed strong interest in such a multimodal AI-powered system. They pointed out several opportunities for these technologies to address their challenges and shared their expectations.
Continuous Monitoring of Symptoms
One of the challenges discussed by clinicians, as highlighted in Section 3.2.1, is the lack of effective ways to monitor symptoms when patients are outside clinical settings. Participants emphasized the potential of wearable sensors to provide continuous, real-time monitoring of patients’ vital signs, such as heart rate and blood pressure. P4 noted the value of such monitoring: “We may be able to find more cardiac patients or cardiac adverse events related to our medications by doing this kind of monitoring.”.
Participants reported that situational factors could arise when patients are in clinical settings, such as “white coat hypertension”, where patients may show elevated blood pressure due to anxiety in a clinical setting. In addition, participants pointed to the advantage of using wearables for patients who are geographically distant from the clinic or unable to visit regularly as they provide remote access to vital health data. P1 highlighted how wearables, particularly in non-clinical setting, could help clinicians gain a more accurate picture of a patient’s condition:
Oftentimes, our patients, especially on the oncology office, they have white code hypertension. So they come and see us, and their blood pressures are sky high, but it’s because they’re nervous. I don’t know what they are like outside of the clinic … Those things can sometimes be helpful. It’s almost like people who have cardiac monitors. ”(P1)
@@
In addition to the potential of wearables in capturing physiological data, participants also saw great potential in using LLM-based CAs to track patient-reported symptoms over time, without requiring manual input from patients. Several participants mentioned that this could ease the burden on patients, many of whom struggle to remember or document their symptoms accurately between appointments. P2 further emphasized the convenience of voice interaction, especially for patients who may have difficulty with traditional reporting methods, such as filling out forms, “This is a really good idea. And that way they talk, and they don’t need to be typing or spending time.
Furthermore, clinicians pointed out that combining wearables and LLM-based CAs could be particularly useful, as each technology addresses different aspects of patient monitoring. These two modalities complement each other and offer the ability to link physiological vital signs with patient-reported symptoms, as P1 commented, “being able to correlate the time with their biometrics could be really helpful.”
AI-based Summarization and Predictive Risk Scores
@@
Several participants noted that navigating extensive medical records could be time-consuming and challenging, and AI-powered summarization tools can help alleviate this burden by providing concise overviews of patient records. P4 echoed this sentiment by highlighting the growing issue of clinician burnout, much of which is driven by the need for documentation in EHRs:
There’s a lot of burnout in medicine, and the burnout is from all of this documentation that we do. I see a huge role of AI in that documentation… fulfill out the EHR data automatically.”(P4)
@@
Also, participants saw promise in AI-generated predictive tools, particularly for assessing cardiotoxicity risk. As mentioned in Section  3.2, clinicians currently rely on a limited set of risk factors to determine which patients may be at higher risk. Participants emphasized the lack of standardized tools for predicting cardiotoxicity and expressed their need for such tools: “I have not come across any risk scoring for cardiotoxicity. So I think that would be helpful.” Participants also noted that AI-driven risk scores could be dynamic and adjust as patient data are updated, offering them a continuous risk assessment. For example, P7 described how AI could assist in this ongoing process:
@@
In addition to its potential, clinicians stressed the importance of the explainability of AI models. They expressed the need for AI models to clearly communicate what risk scores mean and what actions should follow. Without explanations, they may be unsure how to act on the AI’s predictions. For instance, P1 discussed the kind of explanation needed: “like the trigger event of having this higher score is maybe consideration for cardiology referral.”
The need for explainability is also closely tied to how clinicians perceive AI in their workflows. For many participants, they expressed the same sentiment that AI tools should serve as assistants, rather than replace, their clinical judgements. This view reinforces the need for explainability that clinicians want to retain control over final decisions and have AI to provide actionable insights in a way that is understandable, allowing them to comprehend how AI arrived at specific recommendations or risk scores.
3.2.3. Design Revision Suggestions
Participants provided critical feedback that reflected their expectations and highlighted several areas for improvement, allowing us to refine the system and UI design to better align with clinicians’ decision-making workflow.
Monitoring Key Clinical Symptoms and Metrics.
Clinicians emphasized the importance of continuously monitoring a comprehensive set of physiological parameters and symptoms that are clinically important to cardiotoxicity. Many clinicians mentioned heart rate, blood pressure, oxygen saturation, and, where possible, additional indicators such as blood sugar levels. They were curious about the potential of wearable devices to track these diverse metrics, particularly in patients with complex health profiles. Clinicians also noted the importance of monitoring critical symptoms such as chest discomfort, palpitations, shortness of breath, which are typical early indicators of cardiotoxicity.
Additionally, clinicians suggested the need for patient-specific baselines that can alert the care team only when significant deviation occur, rather than bombarding them with excessive data. For example, a sudden increase in the heart rate that deviates from a patient’s established pattern could trigger an alert for further investigation, as P6 suggested, “The care team really should only be alerted if there’s a significant change”
Enhanced Visualizations for Critical Data Interpretation.
Clinicians expressed a clear need for a dynamic and flexible visualization of patient data, with high-level summaries and detailed data views. This dual approach would enable quick assessments while also allowing deeper dives into specific trends and variations when necessary. For instance, a summary graph of daily average heart rates can offer a snapshot of a patient’s cardiovascular status, while the option to view hourly data would provide insight into specific fluctuations and their potential causes. P5 emphasized the importance of trends on isolated data points, stating  “The trend is extremely more important than just the one time”.
This flexibility in visualizing data enables clinicians to identify patterns and correlations that might not be immediately obvious from a single metric. For example, being able to drill down into hourly heart rate data can help clinicians understand the context of a daily average that deviates from the baseline, facilitating more informed decision-making.
Clear Explanation of Technology to Prevent Misunderstanding.
@@ 3.2.3. Design Revision Suggestions
While we presented AI modules and LLM-based voice assistants to clinicians, clinicians expressed their expectations that they should be more explainable to prevent misunderstandings and ensure proper usage. For example, understanding the basis of AI risk scores is crucial for clinicians to trust and effectively use the system. They would like to know how they are calculated and what they present.
@@
For LLM-based voice assistant, they expressed the need to provide clear explanations about how AI responds to different patient scenarios. As P8 pointed out, understanding when to escalate care based on AI alerts is crucial: “Shortness of breath, discomfort, and passing out should always prompt contact with the healthcare provider…something that is acute, that is changing, should be evaluated by a healthcare professional.”

### v3

@@ 3. Participatory Design
Although prior research has identified four key challenges in managing cancer treatment-induced cardiotoxicity and explored the potential of emerging technologies, clinicians’ perspectives on specific design requirements for integrating these technologies into clinical workflows remain underexplored. To bridge this gap, we conducted a participatory design (PD) study with clinical experts specializing in cancer treatment-induced cardiotoxicity, holding sessions individually. The study aims to (1) map how the opportunities offered by emerging technologies can address these challenges and (2) understand the specific needs of clinician technology design to ensure effective integration into their decision-making workflows. By focusing on clinicians’ insights, we sought to align our system design with their practical needs and requirements.
@@
3.1. Study Participants
The PD study aimed to explore how technologies could be tailored to clinicians’ needs, specifically to support their decision-making processes in managing treatment-induced cardiotoxicity. To achieve this, we recruited 11 clinicians with significant experience in this area by convenience sampling (Sedgwick, 2013). Recruitment involved leveraging professional networks, referrals from colleagues, and participant recommendations, ensuring a diverse pool of experts. All participants were affiliated with a North American hospital, representing key specialties – oncology and cardiology – that are central to decision-making for cancer treatment-induced cardiotoxicity. Table 1 presents the demographics of the participants with a randomly assigned number denoted as $P\#$. Participants came from various departments and had a wide range of clinical experience, spanning 1.5 to 13 years.
3.2. Study Procedure
The PD sessions were structured to actively involve clinicians in identifying challenges, exploring technological opportunities, and co-designing solutions to integrate emerging technologies into their workflows. Each individual session was divided into three phases: (1) entry questions, (2) participatory design with an initial User Interface (UI) draft, and (3) exit questions.
We aimed to ensure all participants had a shared conceptual understanding of the technologies explored in RPM and AI-CDSS literature. To achieve this, before PD sessions, we developed an initial low-fidelity UI draft (Figure 2) for a clinician-oriented information dashboard. This prototype served as a probe to facilitate the co-design process, focusing on eliciting participants’ broad feedback. Each feature was represented with placeholders or illustrative examples to demonstrate potential outputs and functionalities, helping bridge gaps in technical literacy and creating a shared foundation for discussion. Rather than presenting a fixed design, we offered multiple design options, varying in both content (e.g., visualizing different types of physiological data such as heart rate and oxygen saturation) and design (e.g., using line charts, bar charts, and dot plots). Examples of such design options are included in Figure 3 as prompts for participant feedback. The initial UI design was informed by prior literature and included conceptual representations of key features such as wearable devices, LLM-VA, and AI-driven prediction models. These conceptual elements were intended to spark discussions around potential functionalities, interactions, and design considerations for integration into clinical workflows.
At the start of each session, participants were asked to recall a recent case involving cancer treatment-induced cardiotoxicity, contextualizing the challenges in monitoring and detection within their workflows. Following this, participants were asked to describe their current use of technologies in these workflows, share their familiarity with emerging tools, and provide their perspectives on potential technologies being studied or proposed for decision-making support in this domain.
During the PD phase, participants actively contributed to the refinement of the initial UI draft. They asked clarifying questions, proposed design revisions, and suggested new features with specific examples. They also envisioned practical use cases for particular scenarios and raised concerns about potential limitations. We concluded each PD session with exit questions to gather final reflections and additional insights from participants. All interview questions can be found in Appendix  A.
Figure 2. Our participatory design session. A participant is suggesting design revisions on the initial UI.
[…]
A screenshot showing the initial user interface prototype for a clinician-facing cardiotoxicity monitoring dashboard, presented during a participatory design session. On the left side of the screen, a Patient List displays the names and details of patients being monitored, including their age and gender. The selected patient, Emily Johnson, has her profile displayed on the right side. This includes her cancer type (Breast Cancer), stage (IIA), and treatment type (Chemotherapy), along with hospitalization dates. The dashboard provides options for viewing Wearable Sensor Data, a Voice Assistant Log, and patient-specific data visualizations. In the center, a graph displays trends in heart rate and oxygen saturation over time, with maximum and minimum values tracked for each metric. Additional buttons allow clinicians to select and view other types of data, such as steps and sleep. On the right side of the screen, a Cardiotoxicity Risk Score of 7/10 is shown, along with options to view Summary and Detailed Data tabs. The summary presents an overall view of the patient’s data and AI model predictions, while the detailed data tab allows clinicians to drill down into specific data for a selected day.
Figure 3. Examples of low-fidelity visualization options for heart rate data used to elicit clinician feedback.
[…]
A set of three side-by-side low-fidelity visualization options for heart rate data designed to elicit clinician feedback during participatory design sessions. The first visualization is a line chart showing trends in maximum and minimum heart rate values over time. The second visualization is a bar chart displaying the same maximum and minimum heart rate values across different time points. The third visualization is a dot chart representing individual heart rate data points, distributed along a timeline. Each chart is labeled ”Wearable Sensor Data” with selectable options for additional metrics, such as ”Oxygen Saturation” and ”Steps.”
@@ 3.2. Study Procedure
The sessions were conducted remotely via Zoom by the first author and lasted 35 to 50 minutes. All sessions were recorded and transcribed with the participants’ consent. Two members of the research team independently analyzed the transcripts using inductive coding and thematic analysis (Braun and Clarke, 2012, 2019). Through iterative discussions, the codes were refined into a consensus codebook to ensure an accurate representation of the data, minimizing redundancy and merging overlapping themes. The transcripts were re-coded based on the finalized codebook (Appendix  2). During the analysis, we observed that thematic saturation was naturally achieved. As we continued coding, no new themes emerged from the data after analyzing the transcripts of the 11 participants.
Our research team combined expertise in HCI, clinical practice (with specialized knowledge in cancer treatment-induced cardiotoxicity), and AI/ML, enabling a comprehensive understanding of our participants. We acknowledge that our interdisciplinary perspectives influenced the design, analysis, and interpretation of the study. By adopting a user-centered approach, we aimed to prioritize the participants’ voices while using our diverse expertise to contextualize their insights within the broader domain.
@@ 3.3.1. Challenges in Clinical Decision-making for Cancer Treatment-Induced Cardiotoxicity
Previous studies (Cardinale et al., 2010; Kalam and Marwick, 2013, 2013; Wittayanukorn et al., 2017; Mascolo et al., 2021; Kamphuis et al., 2019; Pai and Nahata, 2000; Sheppard et al., 2013; Wu et al., 2024) have highlighted four main challenges in decision-making for cancer treatment-induced cardiotoxicity as we described in Section  2.1. While many of the challenges our participants described align with these four, they provide more nuanced insights into them.
Symptoms Not just Subtle and Infrequent, But Sometimes Absent in the Early Stages. Our participants reinforced Challenge 1&2 that symptoms are often subtle and infrequent in the early stages and difficult to detect during routine clinical assessments or outside of clinical settings. They also highlighted a critical nuance: in some cases, symptoms may be entirely absent, even when patients are experiencing severe conditions. For example, P4 shared a case where a patient, completely asymptomatic at home, was found to have a severe arrhythmia solely because he happened to wear a heart monitor, a device not routinely provided to all patients:
@@
Logistical Barriers and Non-Specific Tools Compound Self-Reporting Limitations. Participants reiterated Challenge 3, emphasizing that self-reporting, the primary means through which clinicians learn about patients’ symptoms at home, is fraught with limitations. They also pointed out factors that exacerbate this challenge, including logistical barriers and design-level limitations, which extend beyond the known issue of health literacy and patients’ tendency to downplay symptoms. One issue raised by participants is the presence of logistical barriers, which possibly hinder the effectiveness of self-reporting even when patients are aware of their symptoms. These barriers, such as delayed communication between patients and healthcare providers, are particularly pronounced in resource-constrained settings, leading to discouragement and increased cognitive load on patients, as described by P1:
@@
Participants also revealed the difficulty in recognizing cardiotoxicity-specific symptoms within current self-reporting systems. Many tools currently in use are designed for general purposes rather than tailored to the nuances of cardiotoxicity. This gap leads to missed or underreported cardiac symptoms. For example, P5 emphasized that the surveys they use are “not specific to cardiotoxicity… unless there’s anything on the exam that suggests a cardiac problem.”
@@
Absence of Risk patterns as an Additional Burden to Clinician Workload. Several participants (P1, P2, P7) expressed their frustration with the absence of reliable patterns or correlations between patient characteristics and cardiotoxicity risk, making it challenging to determine which patients require more intensive monitoring. As P1 noted, “We don’t have a good grasp on what features correlate with the likelihood of someone developing cardiotoxicity.” Such uncertainty often results in reactive decision-making, where clinicians have to wait for symptoms to manifest before intervening rather than being able to anticipate and address risks sooner. The uncertainty surrounding risk factors further complicates an already demanding workload Challenge 4, leaving clinicians to manage the additional cognitive and logistical burden of deciding which patients need closer monitoring. Such delays, compounded by late symptom reporting and diagnostic processes, can lead to belated intervention. P2 highlighted this issue by sharing a case in which delays in reporting and testing prolonged treatment decisions: “The symptoms were reported late… everything was delayed, from the echo to the stress test to the EKG.
3.3.2. Opportunities and Suggested Design Revisions
To explore the potential of technologies that can assist clinicians in addressing the identified challenges, we began by asking about their familiarity with wearables, LLM-VA, and AI-based predictive risk scores. We then introduce these concepts alongside an initial UI draft of a clinician-facing dashboard that integrates these modules to facilitate discussion. Participants expressed strong interest in the system’s potential, highlighting key opportunities, sharing their perceptions, and providing suggestions for design improvements. We summarize the key insights below.
Opportunity 1: Continuous Monitoring of Key Clinical Metrics Using Wearables. Participants highlighted the potential of wearables to provide continuous, real-time monitoring of patient’s vital signs, such as heart rate and blood pressure, which could enable identifying subtle patterns and anomalies in patient health. Reflecting on their experiences, clinicians noted the value of such monitoring, with P4 commenting, “I think we can pick up more patients with this. Or we can pick up patients earlier with this… We may be able to find more cardiac patients or cardiac adverse events related to our medications by doing this kind of monitoring.”. Continuous monitoring was also seen as a way to mitigate situational factors in clinical settings, such as “white coat hypertension”, where patients exhibit elevated blood pressure due to anxiety during clinical visits. This opportunity provided by wearables could help clinicians recognize early and subtle symptoms, potentially responding to Challenge 1.
Building on this opportunity, clinicians further emphasized the importance of continuously monitoring a comprehensive set of physiological parameters that are clinically important to cardiotoxicity, which leads to our Design Revision 1: Monitoring Key Clinical Symptoms. Many clinicians mentioned heart rate, blood pressure, oxygen saturation, and, where possible, additional indicators such as blood sugar levels. They were curious about the potential of wearable devices to track these diverse metrics, particularly in patients with complex health profiles. As P1 explained,
What other things could you track with the wearable? Blood sugar would be another one if that would be possible… a lot of our patients with cardiotoxicity also have metabolic syndrome and these blood sugar issues.”(P1)
The ability to track these metrics in real-time continuously allows for the early detection of signs such as arrhythmias or blood pressure fluctuations, which may not be apparent during routine clinical visits.
Opportunity 2: Remote Monitoring for Geographically Distant Patients. Participants highlighted the advantage of using wearables in providing remote access to vital health data for patients who are geographically distant from the clinic or unable to visit regularly. As P1 explained:
A lot of our patients… travel very long distances to come here for treatment… they may call and be having an event where they feel really lightheaded or like they almost passed out… Having that data, especially if they can’t come in right away, is really helpful.”(P1)
This opportunity responds to Challenge 2 by enabling remote, real-time monitoring through wearables, allowing clinicians to track patient health during gaps in clinical interactions and addressing symptoms that might otherwise go unnoticed.
Extending this opportunity, clinicians highlighted the need for patient-specific baselines that can alert the care team only when significant deviation occurs, rather than bombarding them with excessive data. This need informs our Design Revision 2: Patient-Specific Baseline Alerts for Remote Monitoring. For example, a sudden increase in the heart rate that deviates from a patient’s established pattern could trigger an alert for further investigation, as P6 suggested, “the care team really should only be alerted if there’s a significant change. That would be (when) something is wrong, we would need to intervene. Otherwise, it’s just too much data not useful.”
@@ 3.3.2. Opportunities and Suggested Design Revisions
Opportunity 3: LLM-Based Conversational Agent for Symptom Tracking In addition to the potential of wearables in capturing physiological data, participants also saw great potential in using LLM-based CAs to track patient-reported symptoms over time without requiring manual input from patients. Several participants mentioned that this could ease the burden on patients, many of whom struggle to remember or document their symptoms accurately between appointments. P2 emphasized the convenience of voice interaction, especially for patients who may have difficulty with traditional reporting methods, such as filling out forms, “This is a really good idea. And that way they talk, and they don’t need to be typing or spending time.” Furthermore, clinicians pointed out that combining wearables and LLM-based CAs could be particularly useful, as each technology addresses different aspects of patient monitoring. These two modalities complement each other and offer the ability to link physiological vital signs with patient-reported symptoms, as P1 commented, “being able to correlate the time with their biometrics could be really helpful.” This opportunity potentially addresses Challenge 3 by providing automated symptom tracking, reducing reliance on patient memory, and improving the accuracy of symptom reporting through the correlation of physiological and self-reported data.
Building on this opportunity, clinicians first noted the importance of monitoring critical symptoms such as chest discomfort, palpitations, and shortness of breath, which are typical early indicators of cardiotoxicity. As this need is closely related to Design Revision 1, we now extend it to Monitor Key Metrics and Symptoms. Secondly, as clinicians emphasized the importance of correlating symptoms with physiological data, they expressed a clear need for dynamic and flexible visualization of patient information, combining high-level summaries with detailed data views. This leads to our Design Revision 3: Enhanced Visualizations for Critical Data Interpretation. This dual approach would enable quick assessments while also allowing deeper dives into specific trends and variations when necessary. For instance, a summary graph of daily average heart rates can offer a snapshot of a patient’s cardiovascular status. At the same time, the option to view hourly data would provide insight into specific fluctuations and their potential causes. P5 emphasized the importance of trends on isolated data points, stating, “The trend is extremely more important than just the one time.” This flexibility in visualizing data enables clinicians to identify patterns and correlations that might not be immediately obvious from a single metric. For example, being able to drill down into hourly heart rate data can help clinicians understand the context of a daily average that deviates from the baseline, facilitating more informed decision-making.
@@
Opportunity 4: AI-based Summarization and Predictive Risk Scores Several participants noted that navigating extensive medical records could be time-consuming and challenging, echoing Challenge 4, and pointed out that AI-powered summarization tools can potentially help alleviate this burden by providing concise overviews of patient records. P4 echoed this sentiment by highlighting the growing issue of clinician burnout, much of which is driven by the need for documentation in EHRs: “There’s a lot of burnout in medicine, and the burnout is from all of this documentation that we do. I see a huge role of AI in that documentation… fulfill out the EHR data automatically.”
@@
Also, participants saw promise in AI-generated predictive tools, particularly for assessing cardiotoxicity risk. As mentioned in Section  3.3, clinicians currently rely on a limited set of risk factors to determine which patients may be at higher risk. Participants emphasized the lack of standardized tools for predicting cardiotoxicity and expressed their need for such tools: “I have not come across any risk scoring for cardiotoxicity. So I think that would be helpful.” Participants also noted that AI-driven risk scores could be dynamic and adjust as patient data are updated, offering them a continuous risk assessment. For example, P7 described how AI could assist in this ongoing process:
@@
In addition to its potential, clinicians stressed the importance of the explainability of AI models. They expressed the need for AI models to clearly communicate what risk scores mean and what actions should follow. Without explanations, they may be unsure how to act on the AI’s predictions. For instance, P1 discussed the kind of explanation needed: “like the trigger event of having this higher score is maybe consideration for cardiology referral.” The need for explainability is also closely tied to how clinicians perceive AI in their workflows. For many participants, they expressed the same sentiment that AI tools should serve as assistants rather than replace their clinical judgments.
@@
This need informs our Design Revision 4: Clear Explanation of Technology to Prevent Misunderstanding. While we presented AI modules and LLM-based voice assistants to clinicians, clinicians expressed their expectations that they should be more explainable to prevent misunderstandings and ensure proper usage. For example, understanding the basis of AI risk scores is crucial for clinicians to trust and effectively use the system. As P4 commented,
So you have to see what you build it (AI risk score) on… How will you define the score? The problem with scores is that you have to sit down and manually do that. If something does it for us, we’ll be happy to use it. …AI is supposed to be my assistant, not my replacement.”(P4)
@@
They would like to know how these scores are calculated and what they present. Also, they mentioned the need for LLM-based voice assistants to provide clear explanations about how AI responds to different patient scenarios. As P8 pointed out, understanding when to escalate care based on AI alerts is crucial: “Shortness of breath, discomfort, and passing out should always prompt contact with the healthcare provider… something that is acute, that is changing, should be evaluated by a healthcare professional.”